OpenStudy (anonymous):

how do i find the area of a regular 6 sided polygon with a side length of 4.5?????

OpenStudy (anonymous):

chop the hexagon into 6 equilateral triangles that each look exactly the same, with side lengths 4.5. use your work for finding the area of an equilateral triangle with side lengths 4.5 and multiply by six

OpenStudy (anonymous):

Area = (1/2)a*p where "a" is the apothem and "p" is the perimeter(6*4.5). a = the square root of (4.5^2 - 2.25^2) a=3.9 A=(1/2)(3.9)(27) =52.65(approximately)

OpenStudy (anonymous):

wow! thank you so much!